OpenStudy (anonymous):

Gasoline is pouring into a cylindrical tank of radius 3 feet. When the depth of the gasoline is 4 feet, the depth is increasing at 0.2 ft/sec. How fast is the volume of gasoline changing at that instant? Round your answer to three decimal places.

OpenStudy (wolf1728):

I guess the cylindrical tank is oriented this way?

OpenStudy (anonymous):

I think it is with the base on the bottom, not on the sides

OpenStudy (wolf1728):

Well that makes the calculations a LOT simpler

OpenStudy (anonymous):

So I derived the equation in terms of time and I got \[dv/dt=\pi[2rh dr/dt + r^2 dh/dt]\] and because the radius is a constant, the derivative of r is 0 so I emelimated the 2rh in the formula and was left with \[dV/dt=\pi[9 (0.2) dh/dt]\] so is the answer \[dV/dt = 1.8\pi \]?

OpenStudy (wolf1728):

I updated the graphic :

OpenStudy (wolf1728):

I still have the feeling that this problem is for a horizontal cylindrical tank. Is this for a calculus class? If it is then I'd say it's almost certainly horizontal.

OpenStudy (anonymous):

It's vertical

OpenStudy (wolf1728):

Well okay. (That does make the calculations much easier).